A boat is towed on a river so that its speed in still water would be 3 meters per second. If the river flows at the rate of 2 meters per second determine at what inclination to the direction of flow of the river must the boat be headed so that the motion of the boat may be at right angles to the current, if the width of the river is 40 meters, how long will it take for the boat to cross from one bank to another
1
Expert's answer
2019-12-13T10:09:22-0500
"v_b\cos{\theta}=v_w"

"3\cos{\theta}=2"

The inclination to the direction of flow of the river 

"\theta=48.2\degree"

"t=\frac{d}{v_b\sin{\theta}}=\frac{40}{3\sin{48.2\degree}}=17.9\ s"